OpenStudy (anonymous):

Perform the indicated calculations by first expressing all numbers in scientific notation. 0.00003/6,000,000

OpenStudy (anonymous):

I did 3*10^-5/6*10^6=5*10^-11 is this right?

OpenStudy (luigi0210):

Do you know how to convert each into scientific notation?

OpenStudy (luigi0210):

Yup, your conversions are right so d=far

OpenStudy (luigi0210):

*far

OpenStudy (luigi0210):

But I'm not sure about the answer, hold on.

OpenStudy (anonymous):

Thanks and okay I will

OpenStudy (luigi0210):

Okay, so this your equation right? \[\frac{3*10^{-5}}{6*10^6}\] So if I'm right you take them separately right? \[\frac{3}{6}*(10^{-5}*10^{-6})\]

OpenStudy (anonymous):

Yes this iw that was my equation

OpenStudy (anonymous):

I tthink you can choose to do them seperate or togetherk

OpenStudy (anonymous):

Sorry for the typos

OpenStudy (luigi0210):

Yup! You were close on the answer, but you forgot that 3/6=.5 So it'd be: \[.5*10^{-11}\] So just move it over one more time and you got your answer :)

OpenStudy (anonymous):

Oh I see thanks

OpenStudy (luigi0210):

Anytime :)